please answer all
media%2F654%2F654f69ac-e8a8-4a41-924d-f0
media%2Ffb9%2Ffb9d5268-5c24-4598-a82e-94
A car is going with a constant tangential acceleration (i.e. increasing speed) in a circular path. As seen from above, the path is counter-clockwise. It crosses the point 'A' at time 0.00, at which point its speed is known. Find the direction of its centripetal acceleration at some time after it crosses the point A:Write your answer in degrees. Given: Speed when it crosses point 'A' - V-3.6 m/s, tangential acceleration a 3.6 m/s2, radius of path R 176 m, time after crossing point 'A' T-7s Note: the angle should be the direction of the centripetal acceleration, counterclockwise from the positive X-axis. Question 4 1 pts
Answer #1

(1)

Given,

Velocity at point A, v = 3.6 m/s

Tangential acceleration, at = 3.6 m/s2

Radius of circle, R = 176 m

Now,

Angular velocity, \omega 0 = v/R = 3.6/176

                               = 0.021 rad/s

Angular acceleration, \alpha = a/R = 3.6/176

                              = 0.021 rad/s2

After time, t = 7 s

As we know, angle covered after t sec

=> \theta = \omega 0*t + 1/2*\alpha*t2

         = 0.021*7 + (0.5*0.021*7*7)

         = 0.147 + 0.5145 = 0.6615 radians

1 radian = 57.3 degrees

=> \theta = 0.6615*57.3

         = 37.9 degrees

Since direction of centripetal acceleration is radially inwards,

thus, angle made by centripetal acceleration with +x-axis is

=> \theta = 180 + 37.9

         = 217.9 degrees

(2)

Given,

Force applied, F = 99 N

Mass of the block, M = 4.8 kg

mass of the pulley, m = 0.5 kg

radius of the pulley, r = 0.9 m

Displacement of the block, d = 3.6 m

Thus,

moment of inertia of pulley, I = 1/2*m*r2

Now,

let the velocity attained by the block after moving 3.6 m be v

let the angular velocity attained by the pulley after moving 3.6 m be \omega

Since,

=> Work done = change in kinetic energy

=> F*d = 1/2*M*v2 + 1/2*I*\omega2

Since, I = 1/2*m*r2 and \omega = v/r

=> F*d = 1/2*M*v2 + 1/2*(1/2*m*r2)*(v/r)2

=> F*d = 1/2*M*v2 + 1/4*m*v2

=> F*d = (1/2*M + 1/4*m)*v2

=> 99*3.6 = (4.8/2 + 0.5/4)*v2

=> 99*3.6 = (2.4 + 0.125)*v2 = 2.525*v2

=> v2 = (99*3.6)/(2.525) = 141.15

=> v = V14115

         = 11.88 m/s

So,

Kinetic energy of block, K = 1/2*M*v2 = 0.5*4.8*11.88*11.88

                                           = 338.8 J

(3)

mass of ice, m = 20 g

Initial temperature, T = -6 ^{\circ} C

Final temperature, T' = 113 ^{\circ} C

Now,

Total heat required, E = Heat required to raise the temperature of ice from -6 ^{\circ} C to 0 ^{\circ} C

                                         + Latent heat required to convert ice into water

                                            + Heat required to raise the temperature of water from 0 ^{\circ} C to 100 ^{\circ} C

                                                + Latent heat required to convert water into steam

                                                   + Heat required to raise the temperature of steam from 100 ^{\circ} C to 113 ^{\circ} C

=> E = m*cice*(0-(-6)) + m*Lice + m*cwater*(100-0) + m*Lsteam + m*csteam*(113-100)

        = (20*2.090*6) + (20*333) + (20*4.186*100) + (20*2.26*103) + (20*2.010*13)

        = 250.8 + 6660 + 8372 + 45200 + 522.6

        = 61005.4 J

        = 61.01 * 103 J = 61.01 kJ